OpenStudy (anonymous):

Add.simplify if possible. 4w/w^2-64 + w/w-8 =

OpenStudy (anonymous):



OpenStudy (anonymous):

Hi Alisa, lets have a look.

OpenStudy (anonymous):

The first thing you could realize, is that you have a special form of a polynomial on the bottom of the first term, it is in the form: \[a^2-b^2=(a+b)(a-b)\] Or in your case, it is \[(w)^2-(8)^2 = (w-8)(w+8)\]

OpenStudy (anonymous):

Next, do not worry about common denominator as everyone does, all you have to do is use this formula for adding fractions: \[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

In your case a is 4w, b is (w-8)(w+8), c is w, and d is (w-8), so you get: \[\frac{4w}{w^2-64}+\frac{w}{w-8}=\frac{4w}{(w+8)(w-8)}+\frac{w}{w-8}==\frac{4w(w-8)+w(w-8)(w+8)}{(w-8)(w+8)(w-8)}\]

OpenStudy (anonymous):

factor out (w-8) in the numerator, and cancel terms.

OpenStudy (anonymous):

See I got the 4w/(w+8)(w-8)

OpenStudy (anonymous):

I'm confused...

OpenStudy (anonymous):

Hey, sorry about that I had to run off for a bit.

OpenStudy (anonymous):

I was like hey you left me hanging lol

OpenStudy (anonymous):

I came up with 4w/(w+8)(w-8)

OpenStudy (anonymous):

Let me check real quick, also you can check if you are correct by plugging a number in for w (say 1 or something) into your simplified form and the original. They should be the same if you did it correctly

OpenStudy (anonymous):

Lets try that, if you plug in 1 for your answer we get -(4/63)

OpenStudy (anonymous):

If we plug in 1 to the original we get something different (I just typed it into my calculator)

OpenStudy (anonymous):

So you made an error somewhere. why dont you show me your steps and I can help you along the way.

OpenStudy (anonymous):

Ok, well I can try to explain it again. So you figured out that you can factor the w^2-64, this will leave you with \[\frac{4w}{(w-8)(w+8)}+\frac{w}{(w-8)}\]

OpenStudy (anonymous):

ok so you put 4w(w-8)+w(w-8) / (w-8)(w+)(w-8) so I got 4w/(w+8)(w-8)... I just added

OpenStudy (anonymous):

Or is it 4w^2-32/w-8

OpenStudy (anonymous):

Look at what I just put, this is the simple formula you can use for adding fractions. \[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

Ok thanks!!

OpenStudy (anonymous):

In this is case, "a" is equal to 4w right?

OpenStudy (anonymous):

"b" is equal to (w+8)(w-8) right?

OpenStudy (anonymous):

"c" is equal to w, and "d" is equal to (w-8)

OpenStudy (anonymous):

So plugging that into the formula you get \[\frac{4w(w-8) + w(w+8)(w-8)}{(w+8)(w-8)(w-8)}\]

OpenStudy (anonymous):

Now you can factor (w-8) out of the numerator, since it appears in both terms, this gives you:\[\frac{(w-8)[4w+w(w+8)]}{(w+8)(w-8)(w-8)}\]

OpenStudy (anonymous):

Cancel (w-8) from the numerator and denominator

OpenStudy (anonymous):

\[\frac{42 + w(w+8)}{(w+8)(w-8)}\]

OpenStudy (anonymous):

42 should be 4w*

OpenStudy (anonymous):

Then expand what you have left, since there is no more factoring or simplifying that can be done.

OpenStudy (anonymous):

\[\frac{w^2+12w}{w^2-64}\]

OpenStudy (anonymous):

That is as simplified as it can get

OpenStudy (anonymous):

Ok thanks alot! I'll use your method on similiar problems I have

OpenStudy (anonymous):

Cool, yeah I find that it is a lot easier than worrying about the stupid least common denominator etc. It also works on just regular fractions with numbers.